OpenStudy (sonyalee77):

A line passes through the point (–2, 4), and its y-intercept is (0, –6). What is the equation of the line that is perpendicular to the first line and passes through the point (5, –4)?

OpenStudy (sonyalee77):

@ajprincess @AnImEfReaK

OpenStudy (anonymous):

First find the equation of the first line. Use y=mx + c you have c the y intercept. All you need to do it substitute the x coordinate and the y coordinate in the equation.

OpenStudy (kainui):

ANYTIME you see a question about a line, you start here and just start solving. y=mx+b So what's m and what's b of this line in particular that we need to solve for it? Looks like we need to know some stuff about the other line. What is the relationship of a line that's perpendicular to the other line? Are their slopes different? Do their y-intercepts matter?

OpenStudy (anonymous):

@Sonyalee77 can you figure out the equation of the first line?

OpenStudy (sonyalee77):

y=1/5x-3?

OpenStudy (anonymous):

Use the first coordinate (-2, 4)

OpenStudy (anonymous):

c isn't 3 its the y intercept which is (0 , this one) its in the first line of your question.

OpenStudy (anonymous):